# ParallelComp: Parallel Long-Context Compressor for Length Extrapolation

### v1

@@ 3.2 Parallel Attention Bias
In this section, we provide a theoretical framework for understanding Parallel Attention Bias, extending the concept of attention collapse (Dong et al., 2021) to parallel attention mechanisms described in Section 3.1. Specifically, we analyze the sparsity behavior of the local attention matrices computed over parallel chunks and its implications for efficiency and accuracy.
@@
  • Part 1: For any $\epsilon>0$, the sparsity threshold of effective entries in $A^{c}_{\mathfrak{l}}$ increases with $w$. $\epsilon$ is a user-defined threshold that controls sparsity in the attention matrix. With more chunks ($C$), $\epsilon$ affects the balance between retaining information within each chunk and computational efficiency.
  • Part 2: The number of effective entries $k$ in each row of $A^{c}_{\mathfrak{l}}$ is upper-bounded as:
@@
    where $R$ represents the rank of the sparse attention matrix, which determines the effective dimensionality of retained attention entries, and $\delta$ is a probability bound controlling the confidence level of the sparsity constraint.
@@
Proof sketch of Part 1: From the softmax scaling property, the sparsity threshold for effective entries in $A^{c}_{\mathfrak{l}}$ can be bounded as:
@@
This inequality shows that as $w$ increases, the threshold for retaining effective entries becomes stricter, limiting the number of such entries. Proof sketch of Part 2: Rearranging the above inequality, we derive an upper bound on $k$, the number of effective entries:
@@
This implies that the number of effective entries in each row of the attention matrix is given by $w-k$. Proof sketch of Part 3: Substituting the bound of $k$ into the definition of $|\mathcal{S}_{\epsilon}^{(c)}|$, the number of ineffective entries, we find:
@@
Finally, observing that $R=O(\sqrt{\log(w)})$ ensures that the sparsity growth is bounded as $w\to\infty$. Appendix A provides a more detailed proof. ∎
@@ Discussion.
Theorem 3.1 highlights the inevitability of attention collapse in parallel attention mechanisms. Despite dividing the input sequence into $C$ smaller chunks, the effective number of attention entries within each chunk diminishes as the chunk size $w$ increases. Key observations include: i) Each local attention matrix $A^{c}_{\mathfrak{l}}$ exhibits sparsity behavior similar to the global attention matrix, with most entries becoming negligible for large $w$. ii) When a super-long matrix is input into attention in parallel, attention bias inevitably occurs. The attention mechanism repeatedly focuses on a small number of tokens due to its inherent limitations, even with more information provided. Selecting an appropriate sparsity parameter $\epsilon$ can mitigate this issue. iii) Dividing the input into chunks reduces computational overhead while maintaining sparsity within each chunk.
@@ 4 Empirical Study of Parallel Attention Bias
In this section, we investigate the attention sink phenomenon in parallel attention mechanisms and compare its similarities and differences with the regular attention sink phenomenon. Specifically, we explore the following question:
@@
In this part, we provide a detailed analysis of some bias phenomena in parallel attention mechanisms. We observe in Figure 4 that there are relatively more peaks within the contexts.
@@
i) Similar to the blue local attention, the yellow curve shows the U-shaped attention sink phenomenon repeatedly appearing. ii) Parallel attention and local attention both exhibit severe recency bias, but the effect is significantly mitigated in parallel attention compared to local attention. iii) When computing global attention $A{\mathfrak{g}}$, the model suffers from the most severe recency bias, though it is still less pronounced than within $A^{c}_{\mathfrak{l}}$ (blue line). iv) Compared to the classical attention distribution, i.e., the local attention, the peaks of $A{\mathfrak{g}}$ within the chunk are significantly weakened, indicating that global attention can significantly mitigate recency bias. In other words, the parallel attention mechanism itself can mitigate attention bias.
Q3: Can the calibration strategy alleviate attention bias?
@@
i): From Figure 5, we can see that KV cache eviction exacerbates the bias phenomenon. However, parallel KV cache eviction can achieve a more stable distribution. ii): Evicting sink bias tokens in the early layers may exacerbate attention bias, but evicting them in the deeper layers can mitigate this attention bias. iii): Evicting recency bias tokens in the intermediate layers can mitigate attention bias, while evicting recency bias tokens in the deeper layers redistributes the attention scores obtained by the recency bias tokens to the intermediate tokens. iv): Simultaneously evicting sink bias and recency bias tokens can alleviate attention bias in the intermediate layers (Layer 16). v): From Figure 6, by evicting tokens with abnormally high attention scores, the model can mitigate attention bias, which means it can also alleviate the performance loss. We will further validate this in our experiments.
Methods Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg. NtrvQA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TriviaQA SAMSum PCount PRe Lcc RB-P Max Length 84123 24204 17727 20325 19001 20520 60515 34477 16271 13049 26756 21884 32699 17158 37628 58822 30657 Llama2-7B-chat-hf(4k) FullKV(4k) 18.62 19.53 35.49 31.07 26.15 9.91 25.52 20.87 26.28 62.00 82.68 40.86 5.50 10.50 61.04 55.30 33.21 Dynamic-PI 9.69 20.05 33.10 16.40 23.83 3.62 27.83 18.75 16.53 62.00 67.00 40.37 1.58 5.14 55.30 55.49 28.54 NTK-Aware 13.02 14.25 31.51 29.55 30.64 11.83 28.78 16.96 26.30 62.50 74.88 39.35 4.08 4.50 49.74 49.39 30.46 ChunkLlama 22.97 20.52 33.71 28.91 26.14 13.84 14.84 21.62 18.13 62.50 77.15 40.83 2.03 4.00 59.81 54.33 31.33 InfLLM 18.14 22.11 29.86 30.99 30.74 9.41 26.33 20.63 26.18 62.50 84.24 39.92 3.36 6.00 60.15 55.99 32.91 AttenCalibration-NTK 14.05 12.49 32.52 30.61 31.22 12.84 29.72 18.24 24.40 61.50 72.88 39.54 2.33 3.00 48.86 50.36 30.29 Ours 23.20 17.50 37.07 38.67 32.68 20.22 25.00 22.79 25.84 64.00 84.63 40.67 4.00 31.50 59.37 58.53 36.60 Ours-calibration 24.95 19.07 38.16 39.53 32.62 22.64 25.42 22.82 26.01 63.00 85.41 40.36 5.00 32.50 59.04 58.84 37.21 Ours-compression 23.32 16.97 35.25 39.49 32.47 20.17 24.33 21.97 25.68 63.50 84.46 40.81 4.00 31.50 59.43 58.54 36.37 Ours-calibration-compression 24.04 18.39 38.03 39.89 35.38 22.15 24.26 22.46 24.51 63.50 84.83 40.73 4.00 31.50 57.67 58.48 36.86 Llama3-8B-instruct(8k) FullKV(8k) 24.31 38.13 39.69 44.16 35.66 21.00 28.35 23.06 26.96 73.00 90.13 42.46 4.61 68.50 60.46 56.11 42.29 Dynamic-PI 21.71 36.66 38.24 33.70 35.48 14.28 29.41 22.04 25.55 74.50 82.61 42.62 2.33 85.59 58.22 47.16 40.63 NTK-Aware 25.92 37.54 42.23 48.32 36.96 27.51 33.74 24.13 26.35 50.50 88.84 42.53 7.24 95.61 34.84 39.04 41.33 ChunkLlama 25.01 37.39 43.52 49.37 37.56 30.95 17.57 23.51 19.72 76.00 90.38 42.14 4.71 67.95 61.10 52.57 42.47 InfLLM 19.93 43.52 40.58 48.31 35.99 23.25 30.49 21.60 26.53 74.00 90.93 42.30 8.00 74.00 58.98 52.46 43.18 AttenCalibration-NTK 26.54 37.52 41.13 47.56 38.98 26.51 34.21 23.35 25.64 45.50 89.23 42.21 4.81 93.51 36.86 42.82 41.02 Ours 26.67 39.05 42.66 49.58 40.02 33.25 29.10 24.18 26.74 69.00 91.03 42.07 7.81 92.38 58.84 53.54 45.37 Ours-calibration 26.05 42.32 43.81 49.32 40.81 32.86 28.94 24.81 27.42 68.50 91.14 42.33 7.29 93.10 58.72 54.46 45.74 Ours-compression 26.18 36.56 39.72 47.10 34.89 30.10 27.03 23.86 24.52 67.00 89.55 41.20 7.37 92.29 58.51 52.15 43.61 Ours-calibration-compression 26.37 40.18 44.06 47.83 40.77 33.46 26.91 24.35 26.17 69.00 89.74 40.73 6.99 92.47 57.60 54.31 45.06
@@
Llama2-7B-chat-hf(4k) Methods R.PK R.Num R.KV En.MC Math.F Code.Debug Average Max Length 125k 125k 175k 834k 120k 258k 273k FullKV 1.36 1.86 0.4 0.44 17.43 21.57 7.18 Dynamic-PI 0.17 0.00 0.00 7.42 2.00 21.32 5.15 NTK-Aware 2.54 0.00 0.00 3.06 7.71 18.78 5.35 ChunkLlama 12.88 13.22 0.20 0.87 17.14 22.08 11.07 InfLLM 100.00 96.61 2.40 29.80 16.86 22.34 44.67 AttenCalibration-NTK 0.00 0.00 0.00 1.06 5.71 19.24 4.34 Ours 100.00 97.63 20.60 33.62 19.71 25.13 49.45 Ours-calibration 100.00 98.64 30.20 36.24 19.71 31.47 52.71 Ours-compression 97.80 87.96 5.00 35.81 15.86 27.41 44.97 Ours-calibration-compression 97.97 90.14 10.80 35.46 15.86 28.21 46.41 Llama3-8B-instruct(8k) Methods R.PK R.Num R.KV En.MC Math.F Code.Debug Average FullKV 6.10 6.27 4.80 42.79 38.57 22.34 20.15 Dynamic-PI 0.00 0.00 0.00 28.82 29.71 24.62 13.86 NTK-Aware 3.39 8.47 9.40 35.37 39.43 17.77 18.97 ChunkLlama 3.05 9.15 3.60 13.54 34.29 11.42 12.51 AttenCalibration-NTK 4.58 8.47 12.40 34.28 36.57 22.68 19.83 InfLLM 100.00 99.00 5.00 43.70 23.70 22.08 48.91 Ours 100.00 99.83 92.80 54.59 40.00 22.84 68.34 Ours-calibration 100.00 99.49 93.80 56.77 40.00 23.24 68.88 Ours-compression 100.00 99.83 89.20 55.48 40.00 21.32 67.64 Ours-calibration-compression 100.00 99.83 91.00 56.77 40.00 22.20 68.30 Other proprietary models Models R.PK R.Num R.KV En.MC Math.F Code.Debug Average GPT-4 100.00 100.00 89.00 67.25 60.00 37.06 75.55 Kimi-Chat 98.14 95.42 53.60 72.49 12.57 17.14 58.23 Claude-2 97.8 98.14 65.40 62.88 32.29 17.77 62.38 Other open-source models Models R.PK R.Num R.KV En.MC Math.F Code.Debug Average YaRN-Mistral-7B-128k 92.71 56.61 <5 27.95 17.14 60.00 42.82 Yi-6B-200K 100 94.92 <5 36.68 <5 <5 39.85 Yi-34B-200K 100 100 <5 38.43 <5 25.71 44.86 ChatGLM-3-6B-128K 92.2 80.68 <5 10.48 <5 7.71 32.68
@@
5 Experiment
5.1 Experimental Settings
Models, Baselines, and Tasks
We compare our method with existing length extrapolation approaches, including Position Interpolation (PI) (Chen et al., 2023b), NTK-Aware (LocalLLaMA, 2023b), ChunkLlama (An et al., 2024), AttenCalibration (Yu et al., 2024), and InfLLM (Xiao et al., 2024), on LongBench (Bai et al., 2023) and InfiniteBench (Zhang et al., 2024), evaluating them on Llama2-7B-chat-hf (Touvron et al., 2023) and Llama-3-8B-Instruction (AI, 2024). We also compare our method with the following open-source and closed-source models that have been trained on long-context data: ChatGLM-3-6B-128K (GLM et al., 2024), Kimi-Chat (AI, 2023), Yi-6B-200K (01.AI, 2023a), Yi-34B-200K (01.AI, 2023b), Claude-2 (Anthropic, 2023), Yarn-Mistral-7b-128k (Peng et al., 2023), and GPT-4 (Achiam et al., 2023). Since AttenCalibration only calibrates the attention distribution and lacks length extrapolation capability, we implemented NTK-aware technology for it to achieve length extrapolation (AttenCalibration-NTK). We describe our hyperparameters in the Appendix C.

### v2

@@ 3.2 Parallel Attention Bias
In this section, we develop a theoretical framework to understand Parallel Attention Bias, extending the concept of attention collapse (Dong et al., 2021) to parallel attention as described in Section 3.1. We focus on the sparsity behavior of the local attention matrices computed over parallel chunks and examine its impact on both efficiency and accuracy.
@@
  • Part 1: For any $\epsilon>0$, the sparsity threshold of effective entries in $A^{c}_{\mathfrak{l}}$ decreases as $w$ increases. $\epsilon$ represents a user-defined threshold controlling sparsity in the attention matrix. As the number of chunks ($C$) increases, $\epsilon$ governs the trade-off between preserving information within each chunk and computational efficiency.
  • Part 2: The number of effective entries $k$ in each row of $A^{c}_{\mathfrak{l}}$ is upper-bounded by:
@@
    where $R$ is the rank of the sparse attention matrix, influencing the effective dimensionality of retained attention entries, and $\delta$ is a probability bound controlling the confidence level of the sparsity constraint.
@@
Proof sketch of Part 1: By utilizing the exponential decay property of local attention weights (as derived in Theorem A.1), the sparsity threshold for effective entries in $A^{c}_{\mathfrak{l}}$ can be bounded by:
@@
This inequality indicates that as $w$ increases, the threshold for retaining effective entries becomes stricter, thus limiting the number of such entries.
Proof sketch of Part 2: Rearranging the above inequality, we derive an upper bound on $k$, the number of effective entries:
@@
Thus, the number of effective entries in each row of the attention matrix is $w-k$.
Proof sketch of Part 3: Substituting the bound on $k$ into the definition of $|\mathcal{S}_{\epsilon}^{(c)}|$, the number of ineffective entries, we obtain:
@@
Finally, observing that $R=O(\sqrt{\log(w)})$ ensures that the sparsity growth is bounded as $w\to\infty$. A more detailed proof is available in Appendix A. ∎
@@ Discussion.
Theorem 3.1 emphasizes the inevitability of attention collapse in parallel attention. If we fix the sparsity threshold $\epsilon$ and keep the number of chunks $C$ constant, as the input sequence length increases, the effective number of attention entries within each chunk decreases as the chunk size $w$ increases, despite partitioning the input sequence into $C$ chunks. The key insights include: i) Each local attention matrix $A^{c}_{\mathfrak{l}}$ exhibits sparsity behavior akin to the global attention matrix, with most entries becoming negligible for large $w$. ii) When a long sequence is processed in parallel, attention bias becomes unavoidable, with the attention mechanism consistently focusing on a small subset of tokens due to its inherent limitations, even when more information is available. Choosing an appropriate sparsity parameter $\epsilon$ can mitigate this issue. iii) Dividing the input into chunks reduces computational overhead while preserving sparsity within each chunk, leading to an efficient approximation of global attention.
@@ 4 Empirical Study of Parallel Attention Bias
In this section, we investigate the attention sink phenomenon in parallel attention and compare its similarities and differences with the regular attention sink phenomenon. Specifically, we explore the following question:
@@
In this part, we provide a detailed analysis of bias in parallel attention. We observe in Figure 4 that there are relatively more peaks within the contexts compared to the classic attention mechanism.
@@
i) Similar to the blue local attention, the yellow curve shows the U-shaped attention sink repeatedly appearing in global attention. ii) Parallel attention and local attention both exhibit severe recency bias, but the bias is significantly mitigated in parallel attention compared to local attention. iii) When computing global attention $A_{\mathfrak{g}}$, the model suffers from a more severe recency bias compared to the attention sink, though it is still less pronounced than within $A^{c}_{\mathfrak{l}}$ (blue line). iv) Compared to the classical attention distribution, i.e., the local attention, the peaks of $A_{\mathfrak{g}}$ within the chunk are significantly weakened, indicating that global attention can significantly mitigate recency bias. In other words, the parallel attention itself can mitigate attention bias.
Methods Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code AVG NtrvQA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TriviaQA SAMSum PCount PRe Lcc RB-P Max Length 84123 24204 17727 20325 19001 20520 60515 34477 16271 13049 26756 21884 32699 17158 37628 58822 30657 Llama2-7B-chat-hf(4k) FullKV(4k) 18.62 19.53 35.49 31.07 26.15 9.91 25.52 20.87 26.28 62.00 82.68 40.86 5.50 10.50 61.04 55.30 33.21 Dynamic-PI 9.69 20.05 33.10 16.40 23.83 3.62 27.83 18.75 16.53 62.00 67.00 40.37 1.58 5.14 55.30 55.49 28.54 NTK-Aware 13.02 14.25 31.51 29.55 30.64 11.83 28.78 16.96 26.30 62.50 74.88 39.35 4.08 4.50 49.74 49.39 30.46 ChunkLlama 22.97 20.52 33.71 28.91 26.14 13.84 14.84 21.62 18.13 62.50 77.15 40.83 2.03 4.00 59.81 54.33 31.33 InfLLM 18.14 22.11 29.86 30.99 30.74 9.41 26.33 20.63 26.18 62.50 84.24 39.92 3.36 6.00 60.15 55.99 32.91 AttenCalibration-NTK 14.05 12.49 32.52 30.61 31.22 12.84 29.72 18.24 24.40 61.50 72.88 39.54 2.33 3.00 48.86 50.36 30.29 Ours 23.20 17.50 37.07 38.67 32.68 20.22 25.00 22.79 25.84 64.00 84.63 40.67 4.00 31.50 59.37 58.53 36.60 Ours-calibration 24.95 19.07 38.16 39.53 32.62 22.64 25.42 22.82 26.01 63.00 85.41 40.36 5.00 32.50 59.04 58.84 37.21 Ours-compression 23.32 16.97 35.25 39.49 32.47 20.17 24.33 21.97 25.68 63.50 84.46 40.81 4.00 31.50 59.43 58.54 36.37 Ours-calibration-compression 24.04 18.39 38.03 39.89 35.38 22.15 24.26 22.46 24.51 63.50 84.83 40.73 4.00 31.50 57.67 58.48 36.86 Llama3-8B-instruct(8k) FullKV(8k) 24.31 38.13 39.69 44.16 35.66 21.00 28.35 23.06 26.96 73.00 90.13 42.46 4.61 68.50 60.46 56.11 42.29 Dynamic-PI 21.71 36.66 38.24 33.70 35.48 14.28 29.41 22.04 25.55 74.50 82.61 42.62 2.33 85.59 58.22 47.16 40.63 NTK-Aware 25.92 37.54 42.23 48.32 36.96 27.51 33.74 24.13 26.35 50.50 88.84 42.53 7.24 95.61 34.84 39.04 41.33 ChunkLlama 25.01 37.39 43.52 49.37 37.56 30.95 17.57 23.51 19.72 76.00 90.38 42.14 4.71 67.95 61.10 52.57 42.47 InfLLM 19.93 43.52 40.58 48.31 35.99 23.25 30.49 21.60 26.53 74.00 90.93 42.30 8.00 74.00 58.98 52.46 43.18 AttenCalibration-NTK 26.54 37.52 41.13 47.56 38.98 26.51 34.21 23.35 25.64 45.50 89.23 42.21 4.81 93.51 36.86 42.82 41.02 Ours 26.67 39.05 42.66 49.58 40.02 26.23 29.10 24.18 26.74 69.00 91.03 42.07 7.81 92.38 58.84 53.54 44.93 Ours-calibration 26.89 39.46 42.01 49.88 41.41 26.68 29.17 24.55 26.77 72.50 90.53 42.13 8.02 92.75 58.06 53.97 45.21 Ours-compression 26.18 36.56 39.72 47.10 34.89 24.96 27.03 23.86 24.52 67.00 89.55 41.20 7.37 92.29 58.51 52.15 43.31 Ours-calibration-compression 26.46 37.49 41.28 48.28 36.29 26.68 26.79 24.98 25.18 69.00 90.37 40.72 7.34 91.29 57.30 53.97 44.31
Q3: Can the eviction strategy alleviate attention bias?
@@
i): From Figure 5, we can find that KV cache eviction exacerbates the bias. However, parallel KV cache eviction can achieve a more stable distribution. ii): Evicting sink bias tokens in the early layers may exacerbate attention bias, but evicting them in the deeper layers can mitigate this attention bias. iii): Evicting recency bias tokens in the intermediate layers can mitigate attention bias, while evicting recency bias tokens in the deeper layers redistributes the attention scores obtained by the recency bias tokens to the intermediate tokens. iv): Simultaneously evicting sink bias and recency bias tokens can alleviate attention bias in the intermediate layers (Layer 16). v): As shown in Figure 6, evicting tokens with abnormally high attention scores appears to effectively mitigate attention bias within the model. However, the impact of this strategy on task-specific performance remains uncertain. We will investigate this further in our experiments.
5 Experiment
5.1 Experimental Settings
Models, Baselines, and Tasks.
@@ 5.1 Experimental Settings
Llama2-7B-chat-hf(4k) Methods R.PK R.Num R.KV En.MC Math.F Code.Debug Average Max Length 125k 125k 175k 834k 120k 258k 273k FullKV 1.36 1.86 0.4 0.44 17.43 21.57 7.18 Dynamic-PI 0.17 0.00 0.00 7.42 2.00 21.32 5.15 NTK-Aware 2.54 0.00 0.00 3.06 7.71 18.78 5.35 ChunkLlama 12.88 13.22 0.20 0.87 17.14 22.08 11.07 InfLLM 100.00 96.61 2.40 29.80 16.86 22.34 44.67 AttenCalibration-NTK 0.00 0.00 0.00 1.06 5.71 19.24 4.34 Ours 100.00 97.63 20.60 33.62 19.71 25.13 49.45 Ours-calibration 100.00 98.64 22.80 36.24 19.71 30.20 51.27 Ours-compression 97.80 87.96 5.00 35.81 15.86 27.41 44.97 Ours-calibration-compression 97.97 90.14 10.80 35.46 15.86 28.21 46.41 Llama3-8B-instruct(8k) Methods R.PK R.Num R.KV En.MC Math.F Code.Debug Average FullKV 6.10 6.27 4.80 42.79 38.57 22.34 20.15 Dynamic-PI 0.00 0.00 0.00 28.82 29.71 24.62 13.86 NTK-Aware 3.39 8.47 9.40 35.37 39.43 17.77 18.97 ChunkLlama 3.05 9.15 3.60 13.54 34.29 11.42 12.51 AttenCalibration-NTK 4.58 8.47 12.40 34.28 36.57 22.68 19.83 InfLLM 100.00 99.00 5.00 43.70 23.70 22.08 48.91 Ours 100.00 99.83 92.80 54.59 40.00 22.84 68.34 Ours-calibration 100.00 99.49 93.80 56.77 40.00 23.24 68.88 Ours-compression 100.00 99.83 89.20 55.48 40.00 21.32 67.64 Ours-calibration-compression 100.00 99.83 91.00 56.77 40.00 22.20 68.30 Other proprietary models Models R.PK R.Num R.KV En.MC Math.F Code.Debug Average GPT-4 100.00 100.00 89.00 67.25 60.00 37.06 75.55 Kimi-Chat 98.14 95.42 53.60 72.49 12.57 17.14 58.23 Claude-2 97.8 98.14 65.40 62.88 32.29 17.77 62.38 Other open-source models Models R.PK R.Num R.KV En.MC Math.F Code.Debug Average YaRN-Mistral-7B-128k 92.71 56.61 <5 27.95 17.14 60.00 42.82 Yi-6B-200K 100 94.92 <5 36.68 <5 <5 39.85 Yi-34B-200K 100 100 <5 38.43 <5 25.71 44.86 ChatGLM-3-6B-128K 92.2 80.68 <5 10.48 <5 7.71 32.68
@@
We compare our method with existing length extrapolation approaches, including Position Interpolation (PI) (Chen et al., 2023b), NTK-Aware (LocalLLaMA, 2023b), ChunkLlama (An et al., 2024), AttenCalibration (Yu et al., 2024), APE (Yang et al., 2025b), StarAttention (Acharya et al., 2024), and InfLLM (Xiao et al., 2024), on LongBench (Bai et al., 2023) and InfiniteBench (Zhang et al., 2024), evaluating them on Llama2-7B-chat-hf (Touvron et al., 2023), LLaMA3.1 (Grattafiori et al., 2024), Qwen2.5 (Yang et al., 2025a) and Llama-3-8B-Instruction (AI, 2024). We also compare our method with the following open-source and closed-source models trained on long-context data: ChatGLM-3-6B-128K (GLM et al., 2024), Kimi-Chat (AI, 2023), Yi-6B-200K (01.AI, 2023a), Yi-34B-200K (01.AI, 2023b), Claude-2 (Anthropic, 2023), Yarn-Mistral-7b-128k (Peng et al., 2023), and GPT-4 (Achiam et al., 2023). Since AttenCalibration only calibrates the attention distribution and lacks the capability for length extrapolation, we incorporate NTK-aware techniques to enable this functionality, resulting in AttenCalibration-NTK. Details of our hyperparameters are provided in Appendix C.